OpenStudy (trisarahtops):

A particle moves along the x-axis with position function s(t) = e^cos(x). How many times in the interval [0, 2π] is the velocity equal to 0? 1 2 3 More than 3

OpenStudy (baru):

do you know how to differentiate?

OpenStudy (trisarahtops):

I think so yeah

OpenStudy (baru):

s(t) is the function for postition. the derivative of s(t) with respect to time is velocity. i'm sure the question is actually e^cos(t) so can you find V=ds/dt ?

OpenStudy (trisarahtops):

no it is s(t) = ecos(x). So do i start with ln e^(cos x)

OpenStudy (trisarahtops):

*s(t) = e^cos(x)*

OpenStudy (baru):

the question makes sense only if s(t)=e^cos(t)

OpenStudy (trisarahtops):

oh but my problem has x not t....would that matter or can we continue?

OpenStudy (baru):

we'll it matters. buts lets just consider it a mistake in the text book, and change the question to s(t)=e^cos(t)

OpenStudy (baru):

what you should do is differentiate s(t) with respect to t

OpenStudy (trisarahtops):

so I would have -e^cos(x) sin(x)

OpenStudy (baru):

yes. except its not x its t. now we know that the whole expression is zero when ever sin(t) is zero. do you know when sin( ) is zero in the intervall [0,2pi] ?

OpenStudy (trisarahtops):

does this have to do with the unit circle

OpenStudy (baru):

yes

OpenStudy (trisarahtops):

i don't know ;/

OpenStudy (baru):

\(sin(n\pi) \) =0 where n is any integer so in the interval [0,2pi] we have sin(0), sin(pi) and sin (2pi) (n=0,1 and) 3 times (at t=0 , pi and 2pi)

OpenStudy (trisarahtops):

ooohh okay i see. Thank you :D